OpenStudy (hlambach):

Can someone please explain to me how to do 7,8,and 9 on this Proof? Document in comments.

OpenStudy (hlambach):

You don't need to say the answer, just explain how I get the answer, cuz I'm so lost!

OpenStudy (hlambach):

@terenzreignz

terenzreignz (terenzreignz):

Do you use calculus here?

OpenStudy (hlambach):

nope, geometry

terenzreignz (terenzreignz):

Honestly, if I were faced with such a test, I wouldn't be able to answer it quickly like you did :) However, your answers to the previous questions should suffice, you need the equation of the line tangent to the circle and through the point (5,6) You know its slope is -3/4 So you have a slope and a point. The point-slope form of a circle, if you have the slope m, and a point (a,b) goes as follows... \[\Large y = \color{green}m(x-\color{red}a)+\color{blue}b\]

OpenStudy (hlambach):

:) But what about the x-intercept> is it just 5>

OpenStudy (hlambach):

*? *?

terenzreignz (terenzreignz):

First, what's the equation?

OpenStudy (hlambach):

y = mx +b

terenzreignz (terenzreignz):

Yeah, but specifically, what's the equation for the line tangent to the circle at the point (5,6) ?

OpenStudy (hlambach):

uhhhh...

OpenStudy (hlambach):

do you mean the equation of the circle?

OpenStudy (hlambach):

wait nevermind

OpenStudy (hlambach):

i don't know... :(

OpenStudy (hlambach):

(x – 2)^2 + (y – 2)^2 = 5^2 is that it?

terenzreignz (terenzreignz):

NO, the equation of the line tangent to the circle at the point (5,6) You know that its slope is -3/4 and it passes through the point (5,6) So you have a slope and a point. The point-slope form of a circle, if you have the slope m, and a point (a,b) goes as follows... \[\Large y = \color{green}m(x-\color{red}a)+\color{blue}b\]

OpenStudy (hlambach):

so, the equation is y = m(x-a) + b?

OpenStudy (hlambach):

y = -3/4 (x - 5) + 6

terenzreignz (terenzreignz):

That's right :) Now simplify it.

OpenStudy (hlambach):

but how do i find two variables in one equation?

terenzreignz (terenzreignz):

No, I just meant simplify, like distribute that -3/4 and then combine like terms...

OpenStudy (hlambach):

okay

OpenStudy (hlambach):

y = -3/4x + 3 3/4 + 6 y = -3/4 + 9 3/4

OpenStudy (hlambach):

like that?

OpenStudy (hlambach):

*y = -3/4x + 9 3/4

terenzreignz (terenzreignz):

I suppose :) Now to find the x-intercept, set y = 0 and then solve for x :)

OpenStudy (hlambach):

okay: 0 = -3/4x + 9 3/4 -9 3/4 = -3/4x 13 = x

terenzreignz (terenzreignz):

Sorry, I'm a little lost on dividing mixed-number fractions... (how are you doing that?!) Give me a minute...

terenzreignz (terenzreignz):

oh yeah, that seems to work :D

terenzreignz (terenzreignz):

And to get the y-intercept, equivalently, set x = 0, and solve for y.

OpenStudy (hlambach):

haha -9.75 / .75

OpenStudy (hlambach):

*-.75

terenzreignz (terenzreignz):

Okay, fair enough, now, about that y-intercept? :)

OpenStudy (hlambach):

y = -3/4x + 9 3/4 y = 9.75

terenzreignz (terenzreignz):

Anyway, y-intercept should be easy if you have it in y = mx + b form

terenzreignz (terenzreignz):

yeah, that :D Okay, you've got it, yay XD

OpenStudy (hlambach):

:D

OpenStudy (hlambach):

I would give you ten medals if I could! :D

OpenStudy (hlambach):

Thanks so much!

terenzreignz (terenzreignz):

that won't be necessary... besides, you did most of the work >:)

OpenStudy (hlambach):

:)